In class we are learning how to find roots of equations, and I'm doing homework and need some help with it can you help? 3x^2+2x-1
Email is Baseballnoah928@aol.com

Answer by ewatrrr(24785)

3x^2+2x-1= 0
(3x -1)(x+1) = 0, x = -1, 1/3